# Fast, accurate and lightweight sequential simulation-based inference using Gaussian locally linear mappings

### v2

@@ Abstract
Bayesian inference for complex models with an intractable likelihood can be tackled using algorithms performing many calls to computer simulators. These approaches are collectively known as simulation-based inference (SBI). Recent SBI methods have made use of neural networks (NN) to provide approximate, yet expressive constructs for the unavailable likelihood function and the posterior distribution. However, they do not generally achieve an optimal trade-off between accuracy and computational demand. In this work, we propose an alternative that provides both approximations to the likelihood and the posterior distribution, using structured mixtures of probability distributions. Our approach produces accurate posterior inference when compared to state-of-the-art NN-based SBI methods, while exhibiting a much smaller computational footprint. We illustrate our results on several benchmark models from the SBI literature.
@@ 1 Introduction
We propose a novel methodology achieving state-of-the-art Bayesian inference for models with an intractable likelihood function. We illustrate, via several benchmark models, how our methodology returns accurate inference, even in presence of multimodal posteriors, with a small runtime. Moreover, for numerous simulation studies and applications, machine and statistical learning solutions show an ever growing performance, but are not generally evaluated in terms of resource and energy consumption. Although their environmental impact is more and more documented (Schwartz et al., 2020; Strubell et al., 2019; Thompson et al., 2022), the question of resource efficiency and energy consumption when learning model parameters is not as much studied as the quality of the inference itself, due to the multiple aspects to be taken into account and to the lack of clear metrics that limit the algorithms evaluation. When comparing with widely employed inference engines based on neural-networks, our method considerably lower the energy and memory required for the inference task.
@@
We consider simulation-based inference (SBI) for model parameters $\bm{\theta}$ in stochastic modelling. SBI refers to approaches that bypass the need to have a likelihood function $p(\bm{y}|\bm{\theta})$ (for data $\bm{y}$) available in closed form, or one that is cheap to approximate, by instead considering a generative model $f(\bm{\theta})$ or computer simulator. The simulator depends on $\bm{\theta}$ and possibly on other inputs such as covariates and a stream of pseudorandom numbers. Generating a dataset $\bm{y}^{*}$ by running the simulator $\bm{y}^{*}=f(\bm{\theta}^{*})$ at $\bm{\theta}=\bm{\theta}^{*}$ is then equivalent to simulating $\bm{y}^{*}$ from the unavailable likelihood function. In the following this simulation is written as $\bm{y}^{*}\sim p(\bm{y}|\bm{\theta}^{*})$. Established SBI methods include approximate Bayesian computation (ABC, reviewed in Sisson et al., 2018), synthetic likelihoods (Wood, 2010; Price et al., 2018), particle Markov chain Monte Carlo (Andrieu & Roberts, 2009; Andrieu et al., 2010), and several more that we discuss in Section 2. More recently, methods exploiting neural networks (NN) have gained considerable attention. Neural-based estimators have been used to produce approximated likelihoods or posterior distributions, typically using (discrete) normalizing flows (Rezende & Mohamed, 2015; Papamakarios et al., 2021). NN have been used to approximate the likelihood function in implicit models (Papamakarios et al., 2019; Chen et al., 2021), the posterior distribution (Papamakarios & Murray, 2016; Greenberg et al., 2019; Durkan et al., 2020; Chen et al., 2021) or the likelihood and the posterior distribution simultaneously (Wiqvist et al., 2021; Radev et al., 2023a).
@@
These neural-based posterior and likelihood estimation methods have produced exceptionally flexible inference strategies for models with intractable likelihoods. In this work, we propose to investigate more frugal strategies (Evchenko et al., 2021) that can run with limited resources and a much smaller computational footprint, while returning inference of similar quality as neural-based approaches. For SBI, we propose to investigate structured mixtures of probability distributions whose versatility has been widely recognized for a variety of data and tasks, while not requiring excessive design effort or tuning. Their expressivity makes them good candidates to account for complex multivariate models both for the likelihood and posterior distribution. Moreover, mixture models have a much smaller number of hyperparameters compared to NN, and this makes them more amenable to interpretation and efficient learning (see Table 1). More specifically, we design a new simulation-based inference method named Sequential Mixture Posterior and Likelihood Estimation (SeMPLE). SeMPLE learns approximations for both the posterior $p(\bm{\theta}|\bm{y})$ and the likelihood $p(\bm{y}|\bm{\theta})$ using a structured mixture model obtained via the Gaussian Locally Linear Mapping (GLLiM, Deleforge et al., 2014) method. Interestingly, both approximations are jointly returned following a run of an Expectation-Maximization (EM) algorithm within GLLiM. We then use the GLLiM approximate posterior as a well-informed proposal distribution in a tuning-free Metropolis-Hastings sampler. The procedure is embedded in a sequential strategy, which we compare with state-of-the-art NN-based sequential algorithms for posterior inference, namely SNL (Papamakarios et al., 2019) and SNPE-C (Greenberg et al., 2019). Other algorithms could be considered for comparison, however we stick with SNL and SNPE-C and a motivation for this is given in Section 2. Our comparisons are in terms of inference quality as well as resource and energy requirements. In general, we did not compare against ABC samplers such as SMC-ABC, except for the Lotka-Volterra example where such comparison was more suitable, as in Lueckmann et al., 2021 it was shown that standard implementations of ABC approaches usually require a larger number of model simulations, compared to NN-based approaches, although recent research improved the SMC-ABC performance (Picchini & Tamborrino, 2022). We found that, for a given computational budget specified by the total number of model simulations, our method is cheaper to run and produces accurate inference thanks to (i) the informed proposal sampler created by GLLiM and (ii) the computationally faster model training. As an illustration, Figure 1 shows inference for three datasets of the Two Moons model, a model with bimodal posterior that we further analyse in Section 5. The runtime, energy and memory requirements for SeMPLE are considerably lower, see Table 1 and Figure 2. In other examples, (e.g. Figure 3), the speedup provided by SeMPLE is even more dramatic. Overall SeMPLE shows low resources costs that are always of the same magnitude ($<10$ kJ and $<1$ GB), while the requirements for SNL and SNPE-C vary much more with the model dimension or the inference complexity. These results highlight the advantage brought by SeMPLE, which can be run on a minimal configuration while maintaining good performance.
@@
Figure 1: Two Moons model from SBIBM, datasets #1 (top), #8 (middle) and #9 (bottom). Posterior samples after 4 rounds of SeMPLE (column 1), 10 rounds of SNPE-C and SNL (columns 2 and 3), using 10K model simulations. The true posteriors are in column 4.
@@
Figure 2: Two Moons. (left) Median C2ST (the lower the better) and median cumulative runtime in minutes (right) for 10 runs with different data sets vs the number of model simulations. Shaded bands enclose the min and max values.
@@
Table 1: Median resource requirements (across 10 independent repetitions) and number of model hyperparameters $\phi$ for posterior inference with a budget of 10K (two moons and Bernoulli GLM) and 40K (hyperboloid and Ornstein-Uhlenbeck) model simulations. Best (lowest) values are in bold. SNPE-C and SNL used the default SBIBM setup described in Appendix A.
Method CPU+DRAM Peak Memory Usage #Hyperparameters
@@
SNL 66 kJ 39 GB 41,030
SNPE-C 134 kJ 15 GB 102,242
SeMPLE 6.1 kJ 0.93 GB 30,799
Section 2 gives a brief overview of the main SBI approaches. Section 3 describes how GLLiM can be used to construct surrogate posteriors and surrogate likelihoods, while the complete SeMPLE methodology is explained in Section 4. Results from SeMPLE, SNL and SNPE-C are then illustrated in Section 5. Test models, partly reported in the Appendices, include six simulation studies comprising multimodal surfaces (two moons, multiple hyperboloid), stochastic differential equations (Ornstein-Uhlenbeck), generalised linear models (Bernoulli GLM), a Markov jump process with additional measurement error (Lotka-Volterra), and highly correlated posteriors with 20 parameters (twisted-prior). Code is available at https://github.com/henhagg/semple.
@@ 2 Related work
Simulation-based inference, also referred to as likelihood-free inference, is reviewed e.g. in Cranmer et al. (2020) and Pesonen et al. (2023). In a Bayesian framework the focus is on posterior estimation, which in SBI directly relates to the quality of the approximations made to the unavailable likelihood, the posterior distribution or both. This aspect raises the issue of how to evaluate the quality of posterior inference when a reference ground truth posterior is absent by definition. This is still an open question and is further discussed in Section 6. A second important ingredient is sampling efficiency, as simulation-based methods are typically relying on exploration, and on Monte Carlo principles that can be prohibitively expensive. Active or sequential learning is a way to address sample efficiency by gradually discovering and guiding where simulations should be made for a specific given observation. Our proposed method is sequential and we focus on comparisons with NN-based sequential methods for posterior inference, namely SNL (Papamakarios et al., 2019) and SNPE-C (Greenberg et al., 2019). Many other approaches exist for Bayesian inference with intractable likelihoods, such as sequential neural ratio estimation (SNRE, Hermans et al., 2020), the several variants of ABC such as SMC-ABC (Toni et al., 2009; Del Moral et al., 2012), synthetic likelihoods (Wood, 2010; Price et al., 2018) and particle MCMC (Andrieu & Roberts, 2009; Andrieu et al., 2010). Comparisons between several of the above mentioned methods are available in previous works. For example, Greenberg et al. (2019) show that SNPE-C is computationally more efficient than the predecessors [^1] The acronyms SNPE-A and SNPE-B follow the taxonomy in Papamakarios et al. (2019). SNPE-A (Papamakarios & Murray, 2016) and SNPE-B (Lueckmann et al., 2017), where efficient means that, for a given computational budget, SNPE-C approaches the true posterior using a smaller number of model simulations. Lueckmann et al. (2021) show that, for nontrivial case-studies, both SNPE-C and SNRE are more efficient than SNL and SMC-ABC. Moreover, SNPE-C and SNRE perform similarly. The above motivates why in this work we focus on the comparison with SNPE-C using the SBIBM package implementation (Lueckmann et al., 2021) (an exception is the Lotka-Volterra Markov jump process, where we compare against SMC-ABC, as an objective comparison with NN-based methods is difficult here, see section 5.3 for more details). Since SeMPLE includes an MCMC step, same as SNL, it is also relevant to compare against SNL, using its implementation in SBIBM, which is an ameliorated one, compared to the original version of Papamakarios et al. (2019).
@@
Note that there are similarities between our SeMPLE and SNPE-A, as detailed in Appendix C. Briefly, in SNPE-A a $K$ component mixture density network (MDN) is used as a proposal sampler. As such, the MDN means and covariances are parametrised via neural networks. However SNPE-A does not have a built-in strategy to learn each of the $K$ components. Instead, the GLLiM procedure embedded into SeMPLE automatically determines the means and covariance matrices for each of the $K$ components, thus increasing flexibility to the shapes of the posteriors that can be targeted. The fact that SeMPLE does not use MDNs is key to its resource-efficiency, since no complex NN architectures are employed in SeMPLE, and the benefit of the EM algorithm can be exploited towards fitting the mixture models.
@@ 3 Amortized surrogates for likelihoods and posteriors via GLLiM
The inverse regression approach via Gaussian local linear mapping (GLLiM) was first introduced in Deleforge et al. (2014), and has been employed for inference via approximate Bayesian computation (ABC) in Forbes et al. (2022). GLLiM is a parametric mixture model on the joint distribution of $(\bm{y},\bm{\theta})$, denoted by $q_{\tilde{\bm{\phi}}}$ and depending on a set of parameters $\tilde{\bm{\phi}}$. For interpretation or visualization, the model can also be presented as an invertible regression model as follows. In GLLiM the relationship between observations $\bm{y}\in\mathbb{R}^{d}$ and $\bm{\theta}\in\mathbb{R}^{l}$ is assumed to be locally-linear with $K$ components, defined through an additional latent variable $z\in\{1,...,K\}$ as,
$\bm{y}=\sum_{k=1}^{K}\mathbbm{1}\{z=k\}(\tilde{\bm{A}}_{k}\bm{\theta}+\tilde{\bm{b}}_{k}+\tilde{\bm{\epsilon}}_{k}),$ (1)
where $\mathbbm{1}$ denotes the indicator function, and $\tilde{\bm{A}}_{k}\in\mathbb{R}^{d\times l}$ and $\tilde{\bm{b}}_{k}\in\mathbb{R}^{d}$ are matrices and vectors, respectively, defining the affine transformations, while $\tilde{\bm{\epsilon}}_{k}\in\mathbb{R}^{d}$ corresponds to an error term capturing both the observational noise and the modelling error due to assuming an affine approximation for the data. In the following we consider a Gaussian noise, $\tilde{\bm{\epsilon}}_{k}\sim\mathcal{N}_{d}(\bm{0},\tilde{\bm{\Sigma}}_{k}$), and assume $\tilde{\bm{\epsilon}}_{k}$ not depending on $\bm{\theta}$, $\bm{y}$ nor $z$. Therefore, an approximate generative model (while the assumed generative model is denoted $p(\bm{y}|\bm{\theta})$) is given by
@@
Once an estimate for the parameter vector $\tilde{\bm{\phi}}$ (or equivalently $\bm{\phi}$) is available, we have an amortized procedure providing a surrogate likelihood function $q_{\tilde{\bm{\phi}}}(\bm{y}\,|\,\bm{\theta})$, which can be evaluated for any data $\bm{y}$. In particular, if we specify $\bm{y}$ to be a given observed dataset $\bm{y}_{o}$, and plug $\bm{y}_{o}$ into (6), we obtain an approximate (surrogate) likelihood function $q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})$ for our observations $\bm{y}_{o}$. Similarly, an amortized surrogate posterior, $q_{\bm{\phi}}(\bm{\theta}\,|\,\bm{y})$, for a generic $\bm{y}$, can also be derived as a mixture of Gaussian experts distributions from (8). If we plug observed data into (8), then $q_{\bm{\phi}}(\bm{\theta}\,|\,\bm{y}_{o})$ is an approximate posterior based on data. In summary, once $\tilde{\bm{\phi}}$ is estimated via EM, we obtain surrogates of both the likelihood and the posterior in the same GLLiM run. The way this is performed and integrated with the rest of the SeMPLE method is explained in Section 4. Note that it is possible to assume a different distribution for $\tilde{\bm{\epsilon}}_{k}$ than the Gaussian one. For example (1) can be a mixture of generalized Student distributions, resulting in the so-called SLLiM method (Perthame et al., 2018). Both GLLiM and SLLiM are implemented in the R package xLLiM (Perthame et al., 2022) which we employ in our experiments, see Section 5 for details.

### v3

@@ Abstract
Bayesian inference for complex models with an intractable likelihood can be tackled using algorithms performing many calls to computer simulators. These approaches are collectively known as simulation-based inference (SBI). Recent SBI methods have made use of neural networks (NN) to provide approximate, yet expressive constructs for the unavailable likelihood function and the posterior distribution. However, the trade-off between accuracy and computational demand leaves much space for improvement. In this work, we propose an alternative that provides both approximations to the likelihood and the posterior distribution, using structured mixtures of probability distributions. Our approach produces accurate posterior inference when compared to state-of-the-art NN-based SBI methods, even for multimodal posteriors, while exhibiting a much smaller computational footprint. We illustrate our results on several benchmark models from the SBI literature and on a biological model of the translation kinetics after mRNA transfection.
@@ 1 Introduction
We propose a novel methodology achieving state-of-the-art Bayesian inference for models with an intractable likelihood function. We illustrate, via several benchmark models, how our methodology returns accurate inference, even in presence of multimodal posteriors, with a small runtime. Moreover, for numerous simulation studies and applications, machine and statistical learning solutions show an ever growing performance, but are not generally evaluated in terms of resource and energy consumption. Although their environmental impact is more and more documented (Schwartz et al., 2020; Strubell et al., 2019; Thompson et al., 2022), the question of resource efficiency and energy consumption when learning model parameters is not as much studied as the quality of the inference itself, due to the multiple aspects to be taken into account and to the lack of clear metrics that limit the algorithms evaluation. When comparing with widely employed inference engines based on neural-networks, our method considerably lower the energy and memory required for the inference task, while showing competitive inference quality.
@@
We consider simulation-based inference (SBI) for model parameters $\bm{\theta}$ in stochastic modelling. SBI refers to approaches that bypass the need to have a likelihood function $p(\bm{y}|\bm{\theta})$ (for data $\bm{y}$) that is available in closed form, or one that is cheap to approximate, by instead considering a generative model $f(\bm{\theta})$ or computer simulator. The simulator depends on $\bm{\theta}$ and possibly on other inputs such as covariates and a stream of pseudorandom numbers. Generating a dataset $\bm{y}^{*}$ by running the simulator $\bm{y}^{*}=f(\bm{\theta}^{*})$ at $\bm{\theta}=\bm{\theta}^{*}$ is then equivalent to simulating $\bm{y}^{*}$ from the unavailable likelihood function. In the following this simulation is written as $\bm{y}^{*}\sim p(\bm{y}|\bm{\theta}^{*})$. Established SBI methods include approximate Bayesian computation (ABC, reviewed in Sisson et al., 2018), synthetic likelihoods (Wood, 2010; Price et al., 2018), particle Markov chain Monte Carlo (Andrieu & Roberts, 2009; Andrieu et al., 2010), and several more that we discuss in Section 2. More recently, methods exploiting neural networks (NN) have gained considerable attention. Neural-based estimators have been used to produce approximated likelihoods or approximated posterior distributions, typically using (discrete) normalizing flows (Rezende & Mohamed, 2015; Papamakarios et al., 2021). NN have been used to approximate the likelihood function in implicit models (Papamakarios et al., 2019; Chen et al., 2021), the posterior distribution (Papamakarios & Murray, 2016; Greenberg et al., 2019; Durkan et al., 2020; Chen et al., 2021) or the likelihood and the posterior distribution simultaneously (Wiqvist et al., 2021; Radev et al., 2023a).
@@
These neural-based posterior and likelihood estimation methods have produced exceptionally flexible inference strategies for models with intractable likelihoods. In this work, we propose to investigate more frugal strategies (Evchenko et al., 2021) that can run with limited resources and a much smaller computational footprint, while returning inference of similar quality as neural-based approaches. For SBI, we propose to investigate structured mixtures of probability distributions whose versatility has been widely recognized for a variety of data and tasks, while not requiring excessive design effort or tuning. Their expressivity makes them good candidates to account for complex multivariate models both for the likelihood and posterior distribution. Moreover, mixture models have a much smaller number of parameters $\bm{\phi}$ (see Table 1), compared to NN, and this makes them more amenable to interpretation and efficient learning. More specifically, we design a new simulation-based inference method named “Sequential Mixture Posterior and Likelihood Estimation” (SeMPLE). SeMPLE learns approximations for both the posterior $p(\bm{\theta}|\bm{y})$ and the likelihood $p(\bm{y}|\bm{\theta})$ using a structured mixture model obtained via the Gaussian Locally Linear Mapping (GLLiM, Deleforge et al., 2014) method. Interestingly, both approximations are jointly returned, simultaneously, following a run of an Expectation-Maximization (EM) algorithm within GLLiM. We then use the GLLiM approximate posterior as a well-informed proposal distribution in a tuning-free Metropolis-Hastings sampler. The procedure is embedded in a sequential strategy, which we compare with state-of-the-art NN-based sequential algorithms for posterior inference, namely SNL (Papamakarios et al., 2019) and SNPE-C (Greenberg et al., 2019). Other algorithms could be considered for comparison, however we stick with SNL and SNPE-C and a motivation for this is given in Section 2. Our comparisons are in terms of inference quality as well as resource and energy requirements. We did not compare against ABC samplers such as SMC-ABC, except for the Lotka-Volterra example where such comparison was more suitable, as in Lueckmann et al., 2021 it was shown that standard implementations of ABC approaches usually require a larger number of model simulations, compared to NN-based approaches, although recent research improved the SMC-ABC performance (Picchini & Tamborrino, 2024). We found that, for a given computational budget specified by the total number of model simulations, our method is cheaper to run and produces accurate inference thanks to (i) the informed proposal sampler created by SeMPLE and (ii) the fast model training. As an illustration, Figure 1 shows inference for three datasets of the Two Moons model, a model with bimodal posterior that we further analyse in Section 5. The runtime, energy (consumed by the CPU and the DRAM) and memory requirements for SeMPLE are considerably lower, see Table 1, Figure 2 (right) and section 5.5 for details on how energy and memory requirements were measured. In other examples, (e.g. Figure 3), the speedup provided by SeMPLE is even more dramatic (for the model in Supplementary section G.4 we show that SeMPLE is 74 times faster then SNL and 26 times faster than SNPE-C). Overall SeMPLE shows low resources costs that are always of the same magnitude ($<10$ kJ and $<1$ GB), while the requirements for SNL and SNPE-C vary much more with the model dimension or the inference complexity. These results highlight the advantage brought by SeMPLE, which can be run on a minimal configuration while maintaining good performance.
@@
Figure 1: Two Moons model: inference using datasets #1 (top), #2 (middle) and #3 (bottom) from SBIBM. Posterior samples after 4 rounds of SeMPLE (column 1), 10 rounds of SNPE-C and SNL (columns 2 and 3). All methods used a total budget of 10K model simulations. The true posteriors are in column 4.
@@
Figure 2: Two Moons. (left) Median C2ST (the lower the better) and median cumulative runtime in minutes (right) for 10 runs with different data sets vs the number of model simulations. Shaded bands enclose the min and max values. C2ST is the two-samples classifier test, taking values in [0.5,1], see Section 5.
@@
Table 1: Median resource requirements (across 10 independent repetitions) and number of parameters $\phi$ for posterior inference with a budget of 10K (two moons and Bernoulli GLM) and 40K (hyperboloid and Ornstein-Uhlenbeck) model simulations. Best (lowest) values are in bold. SNPE-C and SNL used the default SBIBM setup described in Appendix A.
Method Energy Consumption Peak Memory Usage #parameters $\bm{\phi}$
[kJ] [GB]
@@
SNL 66 39 41,030
SNPE-C 134 15 102,242
SeMPLE 6.1 0.93 30,799
Section 2 gives a brief overview of the main SBI approaches. Section 3 describes how GLLiM can be used to construct surrogate posteriors and surrogate likelihoods, while the complete SeMPLE methodology is explained in Section 4. Results from SeMPLE, SNL and SNPE-C are then illustrated in Section 5. Test models, partly reported in the Appendices, include simulation studies comprising multimodal surfaces (two moons, multiple hyperboloid), stochastic differential equations (Ornstein-Uhlenbeck), generalised linear models (Bernoulli GLM), a Markov jump process with additional measurement error (Lotka-Volterra), highly correlated posteriors with 20 parameters (twisted-prior), and a challenging and partially-observed biological model of the translation kinetics after mRNA transfection. Code is available at https://github.com/henhagg/semple.
@@ 2 Related work
Simulation-based inference, also referred to as likelihood-free inference, is reviewed e.g. in Cranmer et al. (2020) and Pesonen et al. (2023). In a Bayesian framework the focus is on posterior estimation, which in SBI directly relates to the quality of the approximations made to the unavailable likelihood, the posterior distribution or both. This aspect raises the issue of how to evaluate the quality of posterior inference when a reference ground truth posterior is absent by definition. This is still an open question and is further discussed in Section 6. A second important ingredient is sampling efficiency, as simulation-based methods are typically relying on exploration, and on Monte Carlo principles that can be prohibitively expensive. Active or sequential learning is a way to address sample efficiency by gradually discovering and guiding where simulations should be made for a specific given observation. Our proposed method is sequential and we focus on comparisons with NN-based sequential methods for posterior inference, namely SNL (Papamakarios et al., 2019) and SNPE-C (Greenberg et al., 2019). Many other approaches exist for Bayesian inference with intractable likelihoods, such as sequential neural ratio estimation (SNRE, Hermans et al., 2020), the several variants of ABC such as SMC-ABC (Toni et al., 2009; Del Moral et al., 2012), synthetic likelihoods (Wood, 2010; Price et al., 2018) and particle MCMC (Andrieu & Roberts, 2009; Andrieu et al., 2010). Comparisons between several of the above mentioned methods are available in previous works. For example, Greenberg et al. (2019) show that SNPE-C is computationally more efficient than the predecessors [^1] The acronyms SNPE-A and SNPE-B follow the taxonomy in Papamakarios et al. (2019). SNPE-A (Papamakarios & Murray, 2016) and SNPE-B (Lueckmann et al., 2017), where efficient means that, for a given computational budget, SNPE-C approaches the true posterior using a smaller number of model simulations. Lueckmann et al. (2021) show that, for nontrivial case-studies, both SNPE-C and SNRE are more efficient than SNL and SMC-ABC. Moreover, SNPE-C and SNRE perform similarly. The above motivates why in this work we focus on the comparison with SNPE-C using the SBIBM package implementation (Lueckmann et al., 2021) (an exception is the Lotka-Volterra Markov jump process, where we compare against SMC-ABC, as an objective comparison with NN-based methods is difficult here, see section 5.3 for more details). Since SeMPLE includes a Markov chain Monte Carlo (MCMC) step, same as SNL, it is also relevant to compare against SNL, using its implementation in SBIBM, which is an ameliorated one, compared to the original version of Papamakarios et al. (2019).
@@
Note that there are crucial differences between our SeMPLE and SNPE-A, as detailed in Appendix C. Briefly, in SNPE-A a $K$ component mixture density network (MDN) is used as a proposal sampler. As such, the MDN means and covariances are parametrised via neural networks. However SNPE-A does not have a built-in strategy to learn each of the $K$ components. Instead, the GLLiM procedure embedded into SeMPLE automatically determines the means and covariance matrices for each of the $K$ components, thus increasing flexibility to the shapes of the posteriors that can be targeted (see also Wang et al., 2018, where MDNs are employed as proposal functions within a Gibbs sampler). The fact that SeMPLE does not use MDNs, and more generally no NN architectures are employed in SeMPLE [^2] While NN could potentially be incorporated into SeMPLE for the purpose of obtaining summary statistics of the data, as done in eg ABC, this is a separate problem that is not considered in the present work., it results in an easy to train framework, with fewer parameters to learn (see Table 1). Other architectures using NN within proposal samplers are e.g. the neural proposal of Stites et al. (2021) and the mixture of Gaussian copulas of Chen & Gutmann (2019).
@@ 3 Amortized surrogates for likelihoods and posteriors via GLLiM
The inverse regression approach via Gaussian local linear mapping (GLLiM) was first introduced in Deleforge et al. (2014). GLLiM is a parametric mixture model on the joint distribution of $(\bm{y},\bm{\theta})$, denoted by $q_{\tilde{\bm{\phi}}}$ and depending on a set of parameters $\tilde{\bm{\phi}}$. For interpretation or visualization, the model can also be presented as an invertible regression model as follows. In GLLiM the relationship between observations $\bm{y}\in\mathbb{R}^{d}$ and $\bm{\theta}\in\mathbb{R}^{l}$ is assumed to be locally-linear with $K$ components, defined through an additional latent variable $z\in\{1,...,K\}$ as,
$\bm{y}=\sum_{k=1}^{K}\mathbbm{1}_{\{z=k\}}(\tilde{\bm{A}}_{k}\bm{\theta}+\tilde{\bm{b}}_{k}+\tilde{\bm{\epsilon}}_{k}),$ (1)
where $\mathbbm{1}$ denotes the indicator function, and $\tilde{\bm{A}}_{k}\in\mathbb{R}^{d\times l}$ and $\tilde{\bm{b}}_{k}\in\mathbb{R}^{d}$ are matrices and vectors, respectively, defining the affine transformations, while $\tilde{\bm{\epsilon}}_{k}\in\mathbb{R}^{d}$ corresponds to an error term capturing both the observational noise and the modelling error due to assuming an affine approximation for the data. In the following we consider Gaussian noise, $\tilde{\bm{\epsilon}}_{k}\sim\mathcal{N}_{d}(\bm{0},\tilde{\bm{\Sigma}}_{k}$), and assume $\tilde{\bm{\epsilon}}_{k}$ to not depend on $\bm{\theta}$, $\bm{y}$ nor $z$. Therefore, conditionally on component $z=k$, an approximate generative model (while the assumed generative model is denoted $p(\bm{y}|\bm{\theta})$) is given by
@@
Once an estimate for the parameter vector $\tilde{\bm{\phi}}$ (or equivalently $\bm{\phi}$) is available, we have an amortized procedure providing a surrogate likelihood function $q_{\tilde{\bm{\phi}}}(\bm{y}\,|\,\bm{\theta})$, which can be evaluated for any data $\bm{y}$. In particular, if we specify $\bm{y}$ to be a given observed dataset $\bm{y}_{o}$, and plug $\bm{y}_{o}$ into (6), we obtain an approximate (surrogate) likelihood function $q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})$ for observation $\bm{y}_{o}$. Similarly, and using the same training data, an amortized surrogate posterior $q_{\bm{\phi}}(\bm{\theta}\,|\,\bm{y})$, for a generic $\bm{y}$, can also be derived as a mixture of Gaussian experts distributions from (8). If we plug $\bm{y}_{o}$ into (8), then $q_{\bm{\phi}}(\bm{\theta}\,|\,\bm{y}_{o})$ is an approximate posterior based on observed data. In summary, once $\tilde{\bm{\phi}}$ is estimated via EM, we obtain surrogates of both the likelihood and the posterior in the same GLLiM run, simultaneously and using the same training data. The way this is performed and integrated with the rest of the SeMPLE method is explained in Section 4. Note that it is possible to assume a different distribution for $\tilde{\bm{\epsilon}}_{k}$ than the Gaussian one. For example (1) can be a mixture of generalized Student distributions, resulting in the so-called SLLiM method (Perthame et al., 2018). Both GLLiM and SLLiM are implemented in the R package xLLiM (Perthame et al., 2022) which we employ in our experiments, see Section 5 for details. Notice that GLLiM has already been used in SBI, see Forbes et al. (2022), namely they produced an amortized sampler whose posterior draws were then “refined” via a single ABC step. However, in Forbes et al. (2022) only surrogate posteriors were investigated and, in addition, they did not sequentially refine the surrogate posterior and surrogate likelihood. In contrast, our SeMPLE approach makes use of both posterior and likelihood surrogates, to then provide an automatically “tuned” proposal function for MCMC (as described in Section 4) but also to provide a surrogate likelihood, which finds excellent use in the MCMC we construct in Section 4. Then, SeMPLE sequentially refines, in multiple rounds, both the surrogate likelihood and the surrogate posterior (and hence the proposal function), without requiring an ABC step, and hence without the critical tuning/setup decisions that ABC inference entails (eg the choice of distances, threshold values, the determination of suitable proposal functions), and without the large number of simulations that ABC requires.
Another approach that, similarly to GLLiM, could be considered for closed-form approximations of densities without involving neural networks, is the “elliptical process” of Bånkestad et al. (2020), in particular when using their piecewise constant mixing distribution.